one third of a number is 15 less than half of the samwe number what is the number?
Answer by kalants(7)

Your equation is as follows:
(1/3)x = (1/2)x - 15
Subtract (1/2)x from both sides:
-(1/6)x = -15
Multiply both sides by -6 to isolate x:
x = 90